OpenStudy (anonymous):

Just another cute problem: Suppose \(xyz\) is a three digit number such that \(xzy + yxz+yzx+zxy+zyx = 3024\), then can you find \(x \times y \times z\)?

Parth (parthkohli):

XYZ = 100x + 10y + z Is it something related to this?

OpenStudy (anonymous):

Could be, my approach is somewhat different.

Parth (parthkohli):

100x + 10z + y + 100y + 10x + z ...... = 3024

Parth (parthkohli):

let me think

Parth (parthkohli):

122x + 212y + 221x = 3024 I can't go any further :?

OpenStudy (cwtan):

need Permutation?

Parth (parthkohli):

Nope

OpenStudy (anonymous):

@cwtan: I used permutation.

OpenStudy (karatechopper):

what is permutation

Parth (parthkohli):

Hmm...I'm not sure which permutation to use here.

OpenStudy (anonymous):

Number of arrangement.

Parth (parthkohli):

Number of arrangements if order matters.

Parth (parthkohli):

As far as I know, you used permutations only to arrange xyz in different ways

OpenStudy (karatechopper):

i think im gettting to it!!

OpenStudy (karatechopper):

wait..whats the permutation formula?

OpenStudy (karatechopper):

ffm... http://www.mathwords.com/p/permutation_formula.htm in this liinik what do the dots mean in the formula

Parth (parthkohli):

Permutations formula: \(\Large \color{Black}{\Rightarrow _nP_r = {n! \over (n - r)!} }\)

OpenStudy (anonymous):

... this is not a straight forward permutation problem.

OpenStudy (karatechopper):

can u give me a hint..

OpenStudy (cwtan):

I like this question when i am unable to solve it......

OpenStudy (anonymous):

is the answer:\[x\times y\times z = 126\]?

OpenStudy (anonymous):

Yes, approach? :)

OpenStudy (anonymous):

very weird, im pretty sure its nowhere near the best way to look at this.

OpenStudy (anonymous):

222 * 18 - 3024 = 972 --- Whose sum of digits = 18 too.

OpenStudy (anonymous):

Note that 3024 is divisible by 9. Also note that if the three digit number xyz leaves a remainder of r when divided by 9, that any permutation of the digits must leave the same remainder. So adding up those 5 permutations will leave 5r as a remainder. Hence we have this equation:\[5r \equiv 0 \mod 9\]the only solution to this is r = 0 , so the number xyz is divisible by 9, which means the sum of its digits is divisible by 9. Now use what siddhantsharan posted above, using the fact that x+y+z can only be 0, 9, 18, or 27. Its guess and check from there.

OpenStudy (anonymous):

Actually the above post by joemath shortens it down to only 18. As x + y + z must be 15 at least For it to be > 3024. And 222*27 - 3024 will obviously not leave a 3 digit no. Hence 18 has to be correct. Without any guess and check.

OpenStudy (anonymous):

Interesting approach can we generalize it?

OpenStudy (anonymous):

Why do you choose divisibility by 9 in the first place?

OpenStudy (anonymous):

because of how divisibility by 9 is sorta tied to the digits of the number. You can find a numbers remainder when you divide by 9 by adding the digits together. it seemed like a lucky break that 3024 was divisible by 9.

OpenStudy (anonymous):

hmm...actually i take that back, the problem will still be doable even if that sum wasnt divisible by 9. Since we are adding 5 numbers, we will always be able to solve the equation:\[5r\equiv k \mod 9\]Since 5 is invertible (mod 9).

OpenStudy (anonymous):

where k is the remainder of the sum after division by 9.

OpenStudy (cwtan):

So the answer is ? 126?

OpenStudy (karatechopper):

i googled and got 126 but at first when i was solving problem i thought i had to find out what x y and z were..